OpenStudy (anonymous):

Evaluate the line integral (x2+y2)dx+2xydy where is the path of the semicircular arc of the circle x2+y2=64 starting at (8,0) and ending at (−8,0) going counterclockwise.

OpenStudy (turingtest):

did you want to check an answer, or are you stuck?

OpenStudy (anonymous):

\[512\int\limits_{0}^{\pi} \cos^2t sint - \sin^3t\] im stuck at how to integrate this.

OpenStudy (turingtest):

use\[u=\cos t\]and\[\sin^2t=1-\cos^2t\]

OpenStudy (anonymous):

so du =-sinx

OpenStudy (turingtest):

how did you get cos^2t*sint for x^2+y^2dx ?

OpenStudy (anonymous):

i keep getting confused but heres where i started \[\int\limits_{0}^{\pi} ((8cost)^2 + (8sint)^2)(-8sint) + 2(8cost)(8sint)(8cost)dt\]

OpenStudy (turingtest):

yes, and that can be written\[\int_0^\pi512(\sin^2t+\cos^2t)(-\sin t)dt+\int_0^\pi512(2\cos^2t\sin t) dt\]

OpenStudy (anonymous):

\[-512\int\limits_{0}^{?\pi} sint dt +1024\int\limits_{0}^{\pi} \cos^2tsint\] which becomes this

OpenStudy (turingtest):

yes, which you can recombine to form this\[512\int_0^\pi2\cos^2t\sin t-\sin tdt\]

OpenStudy (anonymous):

so would you use u=cost and du=-sint ?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

or could you do \[sint(\cos^2t-1)\] which becomes \[\sin^3t\] ?

OpenStudy (turingtest):

and how to you propose to integrate that?

OpenStudy (anonymous):

oh yeah.haha.

OpenStudy (anonymous):

\[-512\int\limits_{1}^{-1}- 2u^2du - du \] ?

OpenStudy (anonymous):

i mean + du

OpenStudy (turingtest):

yeah, looks good

OpenStudy (turingtest):

oh wait, how did 512 get negative?

OpenStudy (anonymous):

im confused because du=-sint and the sint is positive so (-1/-1) would need to be multiplied right?

OpenStudy (turingtest):

yeah, but you already covered that by changing thethe signs og the terms

OpenStudy (turingtest):

however the double du notation is highly unorthodox and looks screwy to me, so I'd avoid it

OpenStudy (turingtest):

\[512\int_0^\pi2\cos^2t\sin t-\sin tdt\]\[u=\cos t\implies du=-\sin tdt\]\[512\int_{1}^{-1}- 2u^2du +\int_1^{-1} -\sin tdt\]

OpenStudy (turingtest):

sorry I messed up the bounds on the last integral, they should be 0 to pi

OpenStudy (anonymous):

and the answer is 1706.66666666667?

OpenStudy (anonymous):

nevermind thats wrong.

OpenStudy (turingtest):

\[512\int_0^\pi2\cos^2t\sin t-\sin tdt\]\[u=\cos t\implies du=-\sin tdt\]\[512\int_{1}^{-1}- 2u^2du +\int_0^{\pi} -\sin tdt\]let me see what I get...

OpenStudy (turingtest):

\[512\left(\left.-\frac23u^3\right|_1^{-1}+\left.\cos t\right|_0^\pi\right)\]\[512(\frac43-2)=512(-\frac23)\]but I always could have made a mistake...

OpenStudy (anonymous):

its right! yay!

OpenStudy (turingtest):

sweet, I just hope you find your (likely algebra-based) mistake

OpenStudy (anonymous):

thats what i got to too! thanks so much!

OpenStudy (turingtest):

welcome!